OpenStudy (minatsukisaya95):

A balloon is being inflated by pumped air at the rate of 2 cm/s. how fast is the diameter of the balloon increasing when the radius is 1/2 cm.

OpenStudy (anonymous):

could u further explain the part on "A balloon is being inflated by pumped air at the rate of 2 cm/s"? I think u can use differentiation to solve it:D

OpenStudy (minatsukisaya95):

yes it use differentiation but i quite do't understand the question isn't it D=2r. when differentiate it will be dD/dt= 2dr/dt but why the question give the value of r

OpenStudy (minatsukisaya95):

2 cm/s is dr/dt

OpenStudy (anonymous):

dD/dt=(dD/dr) x (dr/dt) =1 x2 =2? kind of confused with this question im assuming dD/dr= 2 x 0.5= 1

OpenStudy (minatsukisaya95):

i assume D=2r so the differentiate is dD/dr=2 what i don't get is how to use the r=1/2 given

OpenStudy (anonymous):

at the instance when r=1/2 the D=1cm?

OpenStudy (minatsukisaya95):

how do we explain that?